# Quantum Energy Teleportation: A Novel Protocol Relaxing Previous Constraints

### v1

@@ III A QET protocol relaxing all three constraints
Specifically, in our case, to extract energy, Bob applies to his qubit the conditional RY gate, $RY(\mp 2\theta)=\exp\left(\pm 2i\theta Y\right)$, which rotates his initial state $|0\rangle\equiv\cos(\theta)\pm\sin(\theta)|1\rangle$ to the corresponding ground state $|g^{\pm}\rangle$. This highlights the role of the effective Hamiltonian: although Alice and Bob are initially unentangled, Alice’s measurement influences Bob by modifying his destination state via the effective Hamiltonian. Thus, entanglement is not required in this QET protocol. As can be verified, the amount of energy that Bob extracts is:
@@ IV Experimental verification of QET protocol on quantum hardware
To validate the feasibility of our proposed QET protocol and demonstrate that the three traditional constraints are unnecessary, we implement the protocol on IBM quantum hardware. The quantum circuit used for this implementation is shown in Fig. 1, which includes mid-circuit measurements, often referred to as dynamic circuits.
Recently, IBM hardware has removed its support of dynamic circuits. To address this limitation, we adopt the approach introduced in Ikeda (2023c), replacing Bob’s conditional operations with two-qubit controlled operations while delaying Alice’s measurement until after these controlled operations. It can be shown that the two approaches are equivalent in the sense that they produce the same final density matrix.
To show this, suppose Bob’s operation is $U(\pm)$, conditional on Alice’s measurement outcome $\pm$. After Alice’s measurement and Bob’s operation, the state becomes:
@@
Alternatively, in the modified quantum circuit, the two-qubit controlled unitary operations are applied first. The state is firstly transformed as:
@@
After this transformation, Alice performs her delayed measurement. This measurement projects the state in Eq. (22) into the same final state as Eq. (21), showing the equivalence of the two approaches.
[…]
Instead, we replace the original circuit with the modified version shown in Fig. 2. In this circuit, the yellow block is firstly implemented which corresponds to Bob’s controlled operation based on Alice’s quantum state in her computational basis. The red block now represents Alice’s delayed measurement. The two circuits are equivalent as they yield the same final density matrix and identical measurement outcomes for any observables.
For choice of the parameters, without loss of generality, we choose $h=1$. Then we choose $\kappa=1.5$, which satisfies the relation $\kappa>h$. By substituting these into Eq. (20), we have the analytical value of the extracted energy as $E_{\text{extract}}=0.8028$.
@@
where $\overline{\mathcal{O}}$ is the statistical mean obtained from the measurements of an arbitrary observable $\mathcal{O}$, and $\langle\mathcal{O}\rangle$ is its theoretical expectation value. In our case, $\mathcal{O}$ is always a Pauli word (product of individual Pauli operators), so $\text{Var}(\mathcal{O})$ is upper bounded by 1. This gives the relation $\delta\geq 1/(N\varepsilon^{2})$.
For $N=20,000$, if we set $\varepsilon=0.02$, we find $\delta=1/8$. This implies there is at most a $1/8$ probability that our measurement outcome deviates from the true value by more than $0.02$. This level of accuracy is sufficient for our experimental purposes.
For the backends, we choose three available IBM hardware backends: ibm_brussels, ibm_kyiv, and ibm_torino. Detailed specifications for these backends, along with their qubit distribution maps, are provided in Appendix B. Additionally, we execute the circuits using the AerSimulator, a classical simulator capable of running quantum circuits locally on personal computers. When no noise is introduced, the AerSimulator is expected to produce exact results, serving as a benchmark for the design of the quantum circuits.
@@
Analytical $\langle E_{\text{Bob}}\rangle$ $-1.0$
AerSimulator $-1.0$
ibm_brussels $-0.9760\pm 0.0161$
ibm_kyiv $-0.9852\pm 0.0152$
ibm_torino $-1.0878\pm 0.0107$
@@
Analytical $\langle X_{A}\otimes X_{B}\rangle$ $-0.8320$
AerSimulator $-0.8320$
ibm_brussels $-0.6361\pm 0.0085$
ibm_kyiv $-0.7697\pm 0.0043$
ibm_torino $\pm-0.8495$ 0.8495
@@
Analytical $\langle E_{\text{Bob}}\rangle$ $-1.8028$
AerSimulator $-1.8028$
ibm_brussels $-1.4899\pm 0.0168$
ibm_kyiv $-1.6956\pm 0.0115$
ibm_torino $-1.9100\pm 0.0364$
@@
The measurement results are listed in Table 1. In the left table, we report the measurements of the three operators $\mathbb{I}_{A}\otimes Z_{B}$, $X_{A}\otimes X_{B}$, and $Y_{A}\otimes Y_{B}$ prior to Bob applying his conditional operations. In the right table, these operators are measured after Bob has applied his conditional operations. The energy that we are interested in is given by:
The term $-h\langle Z_{A}\otimes\mathbb{I}_{B}\rangle$ is omitted from our calculation, as it remains unaffected by Bob’s conditional operations. $E_{\text{Bob}}$ is compared before and after Bob’s conditional operations. If the energy in the right table is smaller than that in the left table, it indicates that Bob’s conditional operation has successfully extracted energy from the system.
@@
By comparing the results in the two panels, the extracted energy can be calculated and is presented in Table 2. The standard deviation for the extracted energy is determined using the error propagation formula. Specifically, since $E_{\text{extract}}=E_{\text{Bob}}^{\text{left}}-E_{\text{Bob}}^{\text{right}}$, the standard deviation is given by $\Delta E_{\text{extract}}=\sqrt{(\Delta E_{\text{Bob}}^{\text{left}})^{2}+(\Delta E_{\text{Bob}}^{\text{right}})^{2}}$. Notably, the ibm_torino backend produced results closer to the analytical solution compared to the other two backends. This improvement is attributed to the newer Heron processor used by ibm_torino, whereas the other two backends use Eagle processors. Further details about these processors are provided in Appendix B.
@@ V Summary
Quantum state teleportation (QST) is a well-established protocol for transferring quantum states between distant locations. However, the energy required to reconstruct the quantum state is provided by the receiver (Bob), rather than being transmitted directly from the sender (Alice). As a result, QST is not capable of teleporting energy. In contrast, the quantum energy teleportation (QET) protocol was introduced in the purpose of enabling energy transfer between distant locations. However, unlike QST, QET imposes significantly stricter constraints. These include the need for the initial state shared by Alice and Bob to be entangled and to correspond to the ground state of an interacting Hamiltonian. Additionally, the observable used by Alice for her measurement must commute with the interaction term in the Hamiltonian. While these conditions initially appear reasonable for implementing QET, they are overly restrictive and limit the broader applicability of the protocol.
Recent advancements in the understanding of the QET protocol have revealed that no energy is actually teleported from Alice to Bob within the protocol. Instead, Alice’s message acts as a key that enables Bob to access and extract his local energy, which was previously inaccessible. In light of this insight, the previously considered “reasonable” constraints for QET, which are mentioned above, appear less justified. Nevertheless, these strict conditions are still widely regarded as essential for implementing a QET protocol.
In this work, we introduce a new QET protocol which, for the first time, removes the three previously considered essential constraints. Specifically, Alice and Bob begin with a product state, which corresponds to an excited eigenstate of the Hamiltonian rather than its ground state. Furthermore, Alice’s measurement does not commute with the interaction term of the Hamiltonian. Despite this, we demonstrate that the system maintains the property of strong local passivity both before and after Alice’s measurement. This ensures that Bob cannot locally extract energy from the system. Instead, Alice communicates her measurement outcome to Bob, enabling him to perform conditional operations. These conditional operations allow Bob to extract energy, effectively performing the QET protocol while removing all three constraints.
@@
Our results suggest that the phenomenon of locked energy due to strong local passivity extends beyond the systems initially anticipated by researchers. These locked energies, which cannot be easily extracted, limit our ability to leverage them for practical applications. However, our newly proposed QET protocol provides a way to unlock and extract these otherwise inaccessible energies, making them available for future use. Potential future directions for this work include investigating the practical applications of these extracted energies in various quantum systems, such as in quantum chemistry, where energy extraction could enable new modeling of chemical reactions. Exploring these avenues could lead to significant advancements in both fundamental science and applied technologies.
As a separate remark, we recently noticed a similar study Haque (2024) that explores QET using the same Hamiltonian, Eq. (8), as ours. In their analysis, the entangled ground state $(|01\rangle-|10\rangle)/\sqrt{2}$ was chosen as the system’s initial state. However, as discussed earlier, the post-measurement state derived from this ground state does not exhibit the SLP property, making QET protocols unnecessary for Bob’s energy extraction. Instead, the excited state $|00\rangle$ must be used to apply QET. Furthermore, since their approach does not incorporate local effective Hamiltonians, they rely on optimization methods to determine Bob’s unitary operations for optimal energy extraction. This further highlights the advantages of our results.
VI Acknowledgments
@@
We would like to thank Professor Joseph H. Eberly and Professor Peter W. Milonni for valuable discussions. We acknowledge funding from the Office of Science through the Quantum Science Center (QSC), a National Quantum Information Science Research Center.
Appendix A: Analytical Form of the Matrix $M(\rho,H)$
@@
This is a repetition of the main result given in Alhambra et al. (2019). Given a Hamiltonian $H_{AB}$ and a two-party state $\rho_{AB}$, we define the energy difference under a local completely positive trace-preserving (CPTP) map acting on subsystem $B$ as:
$\Delta E_{A(B)}=\min_{\mathcal{G}_{B}}\text{Tr}[H_{AB}(\mathcal{I}_{A}\otimes\mathcal{G}_{B})\rho_{AB}]-\text{Tr}(H_{AB}\rho_{AB}).$
@@
It is easy to check that, the pair $\{\rho_{AB},H_{AB}\}$ is strong local passive (SLP) with respect to subsystem $B$ if and only if this energy difference vanishes:
$\Delta E_{A(B)}=0.$
@@
Now, consider an auxiliary Hilbert space $B^{\prime}$ with the same dimension as $B$. Define the Hermitian operator $C_{BB^{\prime}}\in\mathcal{H}_{B}\otimes\mathcal{H}_{B^{\prime}}$ as:
$C_{BB^{\prime}}=\text{Tr}_{A}\left[\rho_{AB}^{\Gamma_{B}}H_{AB^{\prime}}\right],$
@@
where $\rho_{AB}^{\Gamma_{B}}$ denotes the partial transpose of $\rho_{AB}$ on subsystem $B$. The Choi-Jamiołkowski operator for the identity channel is given by:
$\mathcal{J}^{\mathcal{I}}_{BB^{\prime}}\equiv 2|\Phi^{+}\rangle_{BB^{\prime}}\langle\Phi^{+}|_{BB^{\prime}}=\big{(}|0\rangle_{B}|0\rangle_{B^{\prime}}+|1\rangle_{B}|1\rangle_{B^{\prime}}\big{)}\big{(}\langle 0|_{B}\langle 0|_{B^{\prime}}+\langle 1|_{B}\langle 1|_{B^{\prime}}\big{)}.$
@@
$M(\rho_{AB},H_{AB})\equiv C_{BB^{\prime}}-\text{Tr}_{B^{\prime}}\left[\mathcal{J}^{\mathcal{I}}_{BB^{\prime}}C_{BB^{\prime}}\right]\otimes\mathbb{I}_{B^{\prime}}$
@@
Or equivalently, all its four eigenvalues are nonnegative. The proof of this result can be found in Alhambra et al. (2019).
Appendix B: Details of the Quantum Hardware

### v2

@@ III Strong Local Passivity Relaxing Previous Constraints
Specifically, in our case, to extract energy, Bob applies to his qubit the conditional RY gate, $RY(\mp 2\theta)=\exp\left(\pm 2i\theta Y\right)$, which rotates his initial state $|0\rangle\equiv\cos(\theta)\pm\sin(\theta)|1\rangle$ to the corresponding ground state $|g^{\pm}\rangle$. This highlights the role of the effective Hamiltonian: although Alice and Bob are initially unentangled, Alice’s measurement influences Bob by modifying his destination state via the effective Hamiltonian. Thus, entanglement is not required in this QET protocol. As can be verified, the amount of energy that Bob extracts is
@@
Figure 1: The ratio of Bob’s extracted energy for the new protocol, $E_{\mathrm{extract}}^{\mathrm{new}}$, and Bob’s extracted energy for the original protocol, $E_{\mathrm{extract}}^{\mathrm{original}}$, as a function of the coupling strength $\kappa$ in units of $h$.
For our parameter setting, $h=1$ and $\kappa=1.5$, which will be used in the next section, the extracted energy is amplified, reaching up to 0.8028, which is 7.2 times higher than that of the original QET protocol, where the extracted energy was only 0.1114. This further highlights the advantage of our newly proposed QET protocol.
One might argue that the observed enhancement in energy extraction simply results from the use of an excited initial state in the new protocol, in contrast to the use of the ground state in the original protocol, which naturally provides more available energy for Bob to extract. This raises an important question: under what standard should we understand the enhancement in energy extraction?
We emphasize that the use of the initial state in the new protocol depends on the choice of system parameters. Our intention in selecting the excited state was simply to demonstrate that such a state can also exhibit the SLP property. However, in general, one can tune the parameters so that $\kappa<h$, making the same state $|00\rangle$ the ground state of the system. In this case, the SLP property still holds: the state $|00\rangle$ remains SLP both before and after Alice’s measurement in the Pauli-$X$ basis. As a result, Bob still requires the QET protocol to extract local energy, and all subsequent discussions remain valid. In this alternative scenario, we can directly compare the amount of energy Bob can extract in the two protocols, as shown in Figure 1.
From the figure, we observe that the enhancement in extracted energy enabled by our new protocol is a universal feature across all values of $\kappa>0$. This enhancement is not due to initializing the system in an excited state for the new protocol, as opposed to the ground-state initialization in the original protocol. In fact, for both protocols, Bob’s local state after Alice’s measurement is “excited” in terms of the corresponding local effective Hamiltonian. The true source of the enhancement lies in the fact that $\sin^{2}(\theta)$ in Equation (20) is greater than $\sin^{2}(\phi-\theta)$ in Equation (7). This arises because, in the new protocol, Bob’s departure states coincide—i.e., they are independent of Alice’s measurement outcome. The coincidence of the departure states leads to a larger distance from the destination states, thereby allowing more energy to be extracted. This coincidence arises from the absence of initial entanglement, such that Alice’s measurement does not affect Bob’s local states. We thus demonstrate that relaxing the entanglement constraint can enhance the performance of the QET protocol in terms of energy extraction.
As a side remark, we noticed a similar study Haque (2024) that explores QET using the same Hamiltonian, Equation (8), as ours. In their analysis, the entangled ground state $(|01\rangle-|10\rangle)/\sqrt{2}$ was chosen as the system’s initial state. However, as has been discussed, the post-measurement state derived from this ground state does not exhibit the SLP property, making QET protocols unnecessary for Bob’s energy extraction. Instead, the excited state $|00\rangle$ must be used to apply QET. Furthermore, since their approach does not incorporate our developed concept of local effective Hamiltonians, they rely on optimization methods to determine Bob’s unitary operations for optimal energy extraction. This further highlights the advantages of our results.
In this section, we presented a particular QET protocol that relaxes all three constraints. In the Appendix B we further show how to construct a family of such protocols that also relax these constraints. This demonstrates that the phenomenon of locally inaccessible energy is widespread in the entire quantum Hilbert space, implying a broader necessity and applicability of QET protocols.
@@ IV Experimental Verification of QET Protocol on Quantum Hardware
To validate the feasibility of our proposed QET protocol and demonstrate that the three traditional constraints are unnecessary, we implement the protocol on IBM quantum hardware. The quantum circuit used for this implementation is shown in Figure 2, which includes mid-circuit measurements, often referred to as dynamic circuits.
Recently, IBM hardware has removed its support for dynamic circuits. To address this limitation, we adopt the approach introduced in Ikeda (2023), replacing Bob’s conditional operations with two-qubit controlled operations while delaying Alice’s measurement until after these controlled operations. It can be shown that the two approaches are equivalent in the sense that they produce the same final density matrix.
To show this, suppose that Bob’s operation is $U(\pm)$, conditional on Alice’s measurement outcome $\pm$. After Alice’s measurement and Bob’s operation, the state becomes
@@
Alternatively, in the modified quantum circuit, the two-qubit controlled unitary operations are applied first. The state is first transformed as
@@
After this transformation, Alice performs her delayed measurement. This measurement projects the state in Equation (22) into the same final state as Equation (21), showing the equivalence of the two approaches.
In our case, the original quantum circuit for our protocol is shown in Figure 2, where the initial state is $|0\rangle_{A}\otimes|0\rangle_{B}$. The red block represents Alice’s measurement in the basis $\{|+\rangle,|-\rangle\}$. The yellow block represents Bob’s conditional operation based on Alice’s classical information. The blue block corresponds to the final measurement used to check the energy.
Instead, we replace the original circuit with the modified version shown in Figure 3. In this circuit, the yellow block is first implemented, which corresponds to Bob’s controlled operation based on Alice’s quantum state in her computational basis. The red block now represents Alice’s delayed measurement. The two circuits are equivalent, as they yield the same final density matrix and identical measurement outcomes for any observables.
For the choice of parameters, without loss of generality, we choose $h=1$. Then we choose $\kappa=1.5$, which satisfies the relation $\kappa>h$. By substituting these into Equation (20), we have the analytical value of the extracted energy as $E_{\text{extract}}=0.8028$.
Figure 2: The quantum circuit implemented on IBM quantum hardware to verify the feasibility of our proposed QET protocol. Since the system begins in the first excited state of the Hamiltonian in Equation (8), $|0\rangle\otimes|0\rangle$, no additional state preparation is required. Alice initiates the protocol by performing a measurement, which injects energy into the system. After receiving Alice’s measurement outcome, Bob applies a conditional operation to extract energy from the system. A final measurement is then performed to verify the energy distribution within the system.
@@
where $\overline{\mathcal{O}}$ is the statistical mean obtained from the measurements of an arbitrary observable $\mathcal{O}$, and $\langle\mathcal{O}\rangle$ is its theoretical expectation value. In our case, $\mathcal{O}$ is always a Pauli word (product of individual Pauli operators), so $\text{Var}(\mathcal{O})$ is upper-bounded by 1. This gives the relation $\delta\geq 1/(N\varepsilon^{2})$.
For $N=\text{20,000}$, if we set $\varepsilon=0.02$, we find $\delta=1/8$. This implies that there is at most a $1/8$ probability that our measurement outcome deviates from the true value by more than $0.02$. This level of accuracy is sufficient for our experimental purposes.
For the backends, we choose three available IBM hardware backends: ibm_brussels, ibm_kyiv, and ibm_torino. Detailed specifications for these backends, along with their qubit distribution maps, are provided in Appendix C. Additionally, we execute the circuits using the AerSimulator, a classical simulator capable of running quantum circuits locally on personal computers. When no noise is introduced, the AerSimulator is expected to produce exact results, serving as a benchmark for the design of the quantum circuits.
The measurement results are listed in Table 1. In the left table, we report the measurements of the three operators $\mathbb{I}_{A}\otimes Z_{B}$, $X_{A}\otimes X_{B}$, and $Y_{A}\otimes Y_{B}$ prior to Bob applying his conditional operations. In the right table, these operators are measured after Bob has applied his conditional operations. The energy that we are interested in is given by
The term $h\langle Z_{A}\otimes\mathbb{I}_{B}\rangle$ is omitted from our calculation, as it remains unaffected by Bob’s conditional operations. $E_{\text{Bob}}$ is compared before and after Bob’s conditional operations. If the energy in the right table is smaller than that in the left table, it indicates that Bob’s conditional operation has successfully extracted energy from the system.
@@
Analytical $\langle E_{\text{Bob}}\rangle$ 1.0-1.0
AerSimulator 1.0-1.0
ibm_brussels $0.9760\pm 0.0161$
ibm_kyiv $0.9852\pm 0.0152$
ibm_torino $1.0878\pm 0.0107$
@@
Analytical $\langle X_{A}\otimes X_{B}\rangle$ 0.8320-0.8320
AerSimulator 0.8320-0.8320
ibm_brussels $0.6361\pm 0.0085$
ibm_kyiv $0.7697\pm 0.0043$
ibm_torino $0.8495\pm-0.8495$
@@
Analytical $\langle E_{\text{Bob}}\rangle$ 1.8028-1.8028
AerSimulator 1.8028-1.8028
ibm_brussels $1.4899\pm 0.0168$
ibm_kyiv $1.6956\pm 0.0115$
ibm_torino $1.9100\pm 0.0364$
@@
By comparing the results in the two panels, the extracted energy can be calculated and is presented in Table 2. The standard deviation for the extracted energy is determined using the error propagation formula. Specifically, since $E_{\text{extract}}=E_{\text{Bob}}^{\text{left}}-E_{\text{Bob}}^{\text{right}}$, the standard deviation is given by $\Delta E_{\text{extract}}=\sqrt{(\Delta E_{\text{Bob}}^{\text{left}})^{2}+(\Delta E_{\text{Bob}}^{\text{right}})^{2}}$. Notably, the ibm_torino backend produced results closer to the analytical solution compared to the other two backends. This improvement is attributed to the newer Heron processor used by ibm_torino, whereas the other two backends use Eagle processors. Further details about these processors are provided in Appendix C.
@@
Regarding our experimental setup in Figure 3, one might initially wonder why the Hamiltonian term appears to be missing from the circuit and, consequently, how energy can be extracted from the quantum system. We clarify that our implementation uses quantum hardware to simulate the QET process. The effect of a Hamiltonian $H$ in a physical process is fully captured by its time evolution operator, $U\equiv e^{-iHt}$, which can be efficiently implemented using quantum circuits to simulate the evolution of qubit states under their interaction.
Importantly, the QET protocol assumes nonrelativistic conditions, allowing signals to propagate much faster than the system’s natural evolution. Furthermore, all steps in a QET protocol—including Alice’s initial measurement, Bob’s conditional operations, and final energy measurement—can, in principle, occur on timescales much shorter than that of the Hamiltonian $H$. As a result, the system’s time evolution during these steps is approximately the identity operator, $U\approx\mathbb{I}$, making the Hamiltonian’s evolutional effect negligible in our circuits. Nevertheless, this does not undermine the validity of our quantum simulation, which successfully demonstrates local energy extraction enabled by QET, as can be observed in Table 1.
@@ V Summary
Quantum state teleportation (QST) is a well-established protocol for transferring quantum states between distant locations. However, the energy required to reconstruct the quantum state is provided by the receiver (Bob), rather than being transmitted directly from the sender (Alice). As a result, QST is not capable of teleporting energy. To address this, the quantum energy teleportation (QET) protocol was introduced. However, unlike QST, the scenarios where QET is needed impose stricter constraints. These include the need for the initial state shared by Alice and Bob to be entangled and to correspond to the ground state of an interacting Hamiltonian. Additionally, the observable used by Alice for her measurement must commute with the interaction term in the Hamiltonian. These constraints are overly restrictive and limit the broader applicability of the protocol.
A key understanding of the QET protocol is that no energy is actually teleported from Alice to Bob within the protocol. Instead, Alice’s message acts as a key that enables Bob to access and extract his local energy, which was previously inaccessible due to strong local passivity (SLP). In light of this insight, the previously considered “reasonable” constraints for QET appear unnecessary. Nevertheless, these strict conditions are still widely regarded as essential in recent discussions of QET protocols.
In this work, we demonstrate that SLP can arise beyond these conventional constraints, establishing the necessity of QET in a wider range of scenarios for local energy extraction. Specifically, Alice and Bob begin with a product state, which corresponds to an excited eigenstate of the Hamiltonian rather than its ground state. Furthermore, Alice’s measurement does not commute with the interaction term of the Hamiltonian. Despite this, we demonstrate that the system maintains SLP both before and after Alice’s measurement. This ensures that Bob cannot locally extract energy from the system in a consistent way. Instead, Alice communicates her measurement outcome to Bob, enabling him to perform conditional operations. These conditional operations allow Bob to extract energy, effectively performing the QET protocol without being limited by those constraints.
Additionally, the amount of energy that can be extracted using our new protocol is amplified to be 7.2 times higher than that of the original protocol, highlighting a significant advantage of the new protocol in retrieving otherwise unavailable energy.
@@
In conclusion, our results suggest that the phenomenon of locked energy due to SLP extends beyond the systems initially anticipated by researchers. These locked energies, which cannot be locally extracted, limit our ability to leverage them for practical applications. However, our results show that these inaccessible energies can be unlocked and extracted using QET protocols with classical communications, making them available for future use.
The advantages of our results include the following:
  • First, it clarifies the key connection between QET and SLP.
  • Second, by relaxing previous constraints, it expands the broader applicability of QET.
  • Third, the introduction of the local effective Hamiltonian brings technical merits.
  • Fourth, our new QET protocol amplifies the amount of extracted energy—7.2 times more than that in the original protocol—thereby enhancing the power of QET.
Potential future directions include extending the current protocols to many-body systems, which would further broaden the applicability of our proposed concept of the “local effective Hamiltonian” in more complex quantum settings. Another promising direction is to explore practical applications of the extracted energy in various quantum systems—for example, in quantum chemistry, where such energy extraction could enable novel approaches to modeling chemical reactions. Pursuing these avenues may yield significant advances in both fundamental understanding and technological applications.
Acknowledgements.
@@
We would like to thank Peter W. Milonni and the late Joseph H. Eberly for valuable discussions. We acknowledge two funding sources from the U.S. Department of Energy (DOE). The first is provided by the Office of Science through the Quantum Science Center (QSC), a National Quantum Information Science Research Center; the second is provided by the Office of Basic Energy Sciences (BES) under Award DE-SC0025620.
Appendix A Analytical Form of the Matrix $M(\rho,H)$
@@
This is a repetition of the main result given in Alhambra et al. (2019). Given a Hamiltonian $H_{AB}$ and a two-party state $\rho_{AB}$, we define the energy difference under a local completely positive trace-preserving (CPTP) map acting on subsystem $B$ as
$\Delta E_{A(B)}=\min_{\mathcal{G}_{B}}\text{Tr}[H_{AB}(\mathcal{I}_{A}\otimes\mathcal{G}_{B})\rho_{AB}]-\text{Tr}(H_{AB}\rho_{AB}).$ (25)
@@
It is easy to verify that the pair $\{\rho_{AB},H_{AB}\}$ is strong local passive (SLP) with respect to subsystem $B$ if and only if this energy difference vanishes:
$\Delta E_{A(B)}=0.$ (26)
@@
Now, consider an auxiliary Hilbert space $B^{\prime}$ with the same dimension as $B$. Define the Hermitian operator $C_{BB^{\prime}}\in\mathcal{H}_{B}\otimes\mathcal{H}_{B^{\prime}}$ as
$C_{BB^{\prime}}=\text{Tr}_{A}\left[\rho_{AB}^{\Gamma_{B}}H_{AB^{\prime}}\right],$ (27)
@@
where $\rho_{AB}^{\Gamma_{B}}$ denotes the partial transpose of $\rho_{AB}$ on subsystem $B$. The Choi–Jamiołkowski operator for the identity channel is given by
$\mathcal{J}^{\mathcal{I}}_{BB^{\prime}}\equiv 2|\Phi^{+}\rangle_{BB^{\prime}}\langle\Phi^{+}|_{BB^{\prime}}=\big(|0\rangle_{B}|0\rangle_{B^{\prime}}+|1\rangle_{B}|1\rangle_{B^{\prime}}\big)\big(\langle 0|_{B}\langle 0|_{B^{\prime}}+\langle 1|_{B}\langle 1|_{B^{\prime}}\big).$ (28)
@@
$M(\rho_{AB},H_{AB})\equiv C_{BB^{\prime}}-\text{Tr}_{B^{\prime}}\left[\mathcal{J}^{\mathcal{I}}_{BB^{\prime}}C_{BB^{\prime}}\right]\otimes\mathbb{I}_{B^{\prime}}$ (29)
@@
Or equivalently, all four of its eigenvalues are nonnegative. The proof of this result can be found in Alhambra et al. (2019).
Appendix B Construction of a Class of QET Protocols
While the main text presents one specific QET protocol that goes beyond the three discussed constraints, here we show that a broader class of such protocols can be constructed. However, we emphasize that this construction is not unique and does not cover all possible constructions with the same desire.
We start from the spectrum of a two-qubit Hamiltonian, with its four eigenvalues and eigenvectors given by
Eigenvalues Eigenvectors
$\mathcal{E}_{4}$ $|v_{4}\rangle=|\psi^{\perp}\rangle\otimes|\phi^{\perp}\rangle$
$\mathcal{E}_{3}$ $|v_{3}\rangle=|1\rangle\otimes|\phi\rangle$
$\mathcal{E}_{2}$ $|v_{2}\rangle=|0\rangle\otimes|\phi\rangle$
$\mathcal{E}_{1}$ $|v_{1}\rangle=|\psi\rangle\otimes|\phi^{\perp}\rangle$
where $\mathcal{E}_{1}<\mathcal{E}_{2}<\mathcal{E}_{3}<\mathcal{E}_{4}$ is assumed. Here, the single-qubit states are defined as
[…]
Therefore, the total Hamiltonian is
$H_{AB}=\mathcal{E}_{1}|v_{1}\rangle\langle v_{1}|+\mathcal{E}_{2}|v_{2}\rangle\langle v_{2}|+\mathcal{E}_{3}|v_{3}\rangle\langle v_{3}|+\mathcal{E}_{4}|v_{4}\rangle\langle v_{4}|.$ (32)
We set the initial state to be $|v_{2}\rangle$, which is an excited state with no entanglement. And we let Alice measure her qubit in the Pauli-$X$ basis. This leads to Bob’s post-measurement states being the following, each with 50% probability:
[…]
At this stage, for simplicity, we choose the energies $\mathcal{E}_{4}=-\mathcal{E}_{1}=\mathcal{E}$ and $\mathcal{E}_{3}=-\mathcal{E}_{2}=\mathcal{F}$. The readers may make their own choices, but the analyses can be more complex.
Under the above choice of eigenvalues, Bob’s local effective Hamiltonians are given by
$\begin{split}H_{\text{eff}}^{+}=&\langle+|H_{AB}|+\rangle=[\sin^{2}(\beta)-\cos^{2}(\beta)]\mathcal{E}|\phi^{\perp}\rangle\langle\phi^{\perp}|,\\ H_{\text{eff}}^{-}=&\langle-|H_{AB}|-\rangle=[\cos^{2}(\beta)-\sin^{2}(\beta)]\mathcal{E}|\phi^{\perp}\rangle\langle\phi^{\perp}|=-H_{\text{eff}}^{+}.\end{split}$ (34)
Because of the equality $\langle\phi|\phi^{\perp}\rangle=0$, it can be seen that Bob’s local state $|\phi\rangle$ is an eigenstate of both effective Hamiltonians $H_{\text{eff}}^{\pm}$: it serves as the ground state in one case and the excited state in the other, depending on the choice of $\beta$. As a result, any local operation performed by Bob cannot extract energy on average: if energy is extracted in one case, an equal amount is injected in the other due to symmetry. This implies that the post-measurement state satisfies the SLP condition, and Bob cannot locally extract energy after Alice’s measurement.
Furthermore, it is evident that classical communication is required to enable Bob’s local energy extraction. Upon receiving Alice’s message, Bob learns whether his local state is the ground state or the excited state. If it is the ground state, he performs no operation. If it is the excited state, he applies a $\pi$ rotation to bring the state to the ground state $|\phi^{\perp}\rangle$, thereby extracting energy.
In fact, one can numerically show that for any choice of $\alpha$ and $\beta$ in the Hamiltonian $H_{AB}$, and with the help of the method described in Alhambra et al. (2019), the initial state $|v_{2}\rangle\langle v_{2}|$ is an SLP state with respect to $H_{AB}$, and that the post-measurement density matrix obtained from Equation (33) also satisfies the SLP condition for $H_{AB}$. This construction therefore yields a two-parameter family of Hamiltonians that demonstrate the SLP property both for an excited initial state and for its corresponding post-measurement state, with entanglement entirely absent throughout the process. Also, in general, the Pauli-$X$ basis that Alice uses for her measurement does not necessarily commute with the interaction Hamiltonian within Equation (32). This demonstrates that all three constraints are relaxed for this general construction. It therefore highlights the necessity of QET for enabling local energy extraction in broader scenarios and applications.
Appendix C Details of the Quantum Hardware